What is the third longest wavelength in the absorption spectrum of hydrogen?

Express your answer to four significant figures and include the appropriate units.
Answer #1

The Rydberg formula is:

\frac{1}{\lambda }=R\left ( \frac{1}{n_{1}^{2}}-\frac{1}{n_{2}^{2}} \right )

R=Rydberg constant

Consider the atom is on the fundamental state when excited: n1=1

The third wavelength will be for n2=4: ~97.2018*10-9 m.